# UI-AGILE: Advancing GUI Agents with Effective Reinforcement Learning and Precise Inference-Time Grounding

### v1

@@ 2 Related Work
The field of GUI agents has evolved rapidly (Wang et al. 2024; Li and Huang 2025). Following early works like CogAgent (Hong et al. 2024) and SeeClick (Cheng et al. 2024), recent studies rely on SFT to operate directly on visual inputs. Show-UI (Lin et al. 2025) innovates on visual processing efficiency. OS-Atlas (Wu et al. 2025), UGround (Gou et al. 2025) and Aria-UI (Yang et al. 2024) propose novel, large-scale pipelines to collect and synthesize millions of GUI agent trajectories, significantly improving model generalization. Aguvis (Xu et al. 2024) introduced a two-stage training process that explicitly uses VLM-generated Chain-of-Thought (CoT) data to teach planning and reasoning. JEDI (Xie et al. 2025) constructs a refusal part by mismatching existing instructions with unrelated screenshots. Standing out in complexity and scale, UI-TARS (Qin et al. 2025) utilizes the largest dataset and the most intricate training pipeline, which involves SFT and DPO on human annotated CoT data to improve performance. This data-intensive scaling has motivated a shift towards more efficient Reinforcement Learning (RL) paradigms, first explored by UI-R1 (Lu et al. 2025) and GUI-R1 (Luo et al. 2025). InfiGUI-R1 (Liu et al. 2025a) employs Spatial Reasoning Distillation and RL to enhance planning and error recovery capabilities. GUI-G1 (Zhou et al. 2025) leverages Hit-based reward and IoU-based reward for improving GUI agents.
@@ 4.1 Implementation Details
Baselines.
We include detaild descriptions for all baselines in Appendix.
@@
We use the trl framework (von Werra et al. 2020) to implement the cropping-based resampling strategy and reward functions. The sampling process is attempted 4 times at most and is bypassed entirely if the bbox’s dimensions exceed the target crop size. Following previous works (Lu et al. 2025; Luo et al. 2025; Liu et al. 2025a; Zhou et al. 2025), we use Qwen2.5-VL [^1] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct, https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct as base model. Tab. 1 provides the default hyperparameters where cropping factor is the width and height ratio of new attempted image and last attempted image.
@@ 4.3 Agent Capability Evaluation
Following the evaluation setting of OS-Atlas (Wu et al. 2025), we use three metrics: action type prediction accuracy (Type), grounding accuracy (GR), and the overall step success rate (SR). Type accuracy measures the exact match for the predicted action (e.g., click or scroll). For GR, a prediction is considered successful if it falls within a 14% screen-width radius of the ground-truth coordinate. The most holistic metric, SR, deems a step successful only if both the action type and all its associated arguments (e.g., coordinates for a click, direction for a scroll, or text for an input) are correct. For the number of test samples on AndroidControl, we follow OS-Atlas and use 7,708 examples for a fair comparison, while some works (e.g., Aguvis (Xu et al. 2024) and UGround (Gou et al. 2025)) randomly sample 500 action steps from the full AndroidControl test set for testing.
The evaluation is conducted under two distinct settings. In the AndroidControl-Low setting, the agent receives a specific, low-level instruction for each step. In contrast, the AndroidControl-High setting provides the agent with a high-level goal, requiring it to infer the correct action for the current step based on the conversation history and previous actions, thereby testing its multi-step reasoning capability.
As presented in Tab. 4, our UI-AGILE-7B model achieves the best performance (SR: 77.6 and 60.6) compared to all other RFT models including UI-R1-E (SR: 71.37 and 35.88), GUI-R1-3B (SR: 64.41 and 46.55) and GUI-R1-7B (66.52 and 51.67). Notably, our smaller UI-AGILE-3B model with SR of 77.4 and 56.9 also surpasses 7B baselines like GUI-R1-7B (SR: 66.5 and 51.7). This remarkable result demonstrates that the improvements gained from our methods are not confined to improving grounding capability but also translate effectively to better decision-making and execution in complex, multi-step agent scenarios.
@@ 4.4 Ablation Study
Figure 3: Ablation study on grounding benchmarks.
Figure 4: Ablation study on AndroidControl benchmark.
@@ 4.6 Analysis of Inference Time
We report the inference time of our decomposed grounding with selection method on the full ScreenSpot-Pro dataset (Li et al. 2025) using the vLLM framework (Kwon et al. 2023) and one 80G A800 GPU card. As a baseline, the standard grounding approach applied to UI-AGILE-7B without our method completes the benchmark in 30 minutes. When applying our method, the decomposed grounding stage takes 35 minutes. The subsequent VLM-based selection stage requires additional 4 minutes. The modest increase in overhead is a practical trade-off for the substantial gain of grounding accuracy brought by our method.

### v2

@@ 2 Related Work
The field of GUI agents has evolved rapidly (Wang et al. 2024; Li and Huang 2025). Following early works like CogAgent (Hong et al. 2024) and SeeClick (Cheng et al. 2024), recent studies rely on SFT to operate directly on visual inputs. Show-UI (Lin et al. 2025) innovates on visual processing efficiency. OS-Atlas (Wu et al. 2025), UGround (Gou et al. 2025) and Aria-UI (Yang et al. 2024) propose novel, large-scale pipelines to collect and synthesize millions of GUI agent trajectories, significantly improving model generalization. Aguvis (Xu et al. 2024) introduced a two-stage training process that explicitly uses VLM-generated Chain-of-Thought (CoT) data to teach planning and reasoning. JEDI (Xie et al. 2025) constructs a refusal part by mismatching existing instructions with unrelated screenshots. Standing out in complexity and scale, UI-TARS (Qin et al. 2025) utilizes the largest dataset and the most intricate training pipeline, which involves SFT and DPO on human annotated CoT data to improve performance. This data-intensive scaling has motivated a shift towards more effective RL paradigms, first explored by UI-R1 (Lu et al. 2025) and GUI-R1 (Luo et al. 2025). InfiGUI-R1 (Liu et al. 2025a) employs Spatial Reasoning Distillation and RL to enhance planning and error recovery capabilities. GUI-G1 (Zhou et al. 2025) leverages Hit-based reward and IoU-based reward for improving GUI agents.
@@ 4.1 Implementation Details
We use the trl framework (von Werra et al. 2020) to implement the cropping-based resampling strategy and reward functions. The sampling process is attempted 4 times at most and is bypassed entirely if the bbox’s dimensions exceed the target crop size. Following prior works (Lu et al. 2025; Luo et al. 2025; Liu et al. 2025a; Zhou et al. 2025), we use Qwen2.5-VL-3B [^1] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B [^2] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct as base models. Tab. 1 provides the default hyperparameters where cropping factor is the width and height ratio of new attempted image and last attempted image.
@@ 4.3 Agent Capability Evaluation
Following the evaluation setting of OS-Atlas (Wu et al. 2025), we use three metrics: action type prediction accuracy (Type), grounding accuracy (GR), and the overall step success rate (SR). Type accuracy measures the exact match for the predicted action (e.g., click or scroll). For GR, a prediction is considered successful if it falls within a 14% screen-width radius of the ground-truth coordinate. The most holistic metric, SR, deems a step successful only if both the action type and all its associated arguments (e.g., coordinates for a click, direction for a scroll, or text for an input) are correct.
Following OS-Atlas, we use 7,708 examples for a fair comparison, while some works (e.g., Aguvis (Xu et al. 2024) and UGround (Gou et al. 2025)) randomly sample 500 action steps for testing. The evaluation is conducted under two distinct settings. In the AndroidControl-Low setting, the agent receives a specific, low-level instruction for each step. In contrast, the AndroidControl-High setting provides the agent with a high-level goal, requiring it to infer the correct action for the current step based on the conversation history and previous actions, thereby testing its multi-step reasoning capability.
As presented in Tab. 4, our UI-AGILE-7B model achieves the best performance (SR: 77.6 and 60.6) compared to other RFT models including UI-R1-E (SR: 71.37 and 35.88), GUI-R1-3B (SR: 64.41 and 46.55) and GUI-R1-7B (SR: 66.52 and 51.67).
Notably, our smaller UI-AGILE-3B model with SR of 77.4 and 56.9 also surpasses 7B baselines like GUI-R1-7B (SR: 66.5 and 51.7). For Type, a similar trend can be observed.
For GR, UI-AGILE archives the best performance in most cases except that GUI-R1-7B shows higher GR (65.6) in the AndroidControl-High setting. We attribute it to different training philosophies: their prolonged training on a smaller dataset for 9 epochs may foster specialization, whereas our training on a larger, more diverse dataset for 2 epochs prioritizes generalization. This hypothesis is supported by our model’s dominant performance on dedicated grounding benchmarks ScreenSpot-Pro and ScreenSpot-v2.
Overall, the remarkable results on AndroidControl demonstrate that the improvements gained from our methods are not confined to improving grounding capability but also translate effectively to better decision-making and execution in complex, multi-step agent scenarios.
@@ 4.4 Ablation Study
Figure 3: Ablation study on grounding benchmarks.
Figure 4: Ablation study on AndroidControl benchmark.
@@ 4.6 Analysis of Inference Time
We report the inference time of our decomposed grounding with selection method on the full ScreenSpot-Pro dataset (Li et al. 2025) using the vLLM framework (Kwon et al. 2023) and one 80G A800 GPU card.
As a baseline, the standard grounding approach applied to UI-AGILE-7B without our method completes the benchmark in 30 minutes. When applying our method, the decomposed grounding stage takes 35 minutes. The subsequent VLM-based selection stage requires additional 4 minutes. The modest increase in overhead is a practical trade-off for the substantial gain of grounding accuracy brought by our method.